Find the literature ksp value for Mn(OH )2 . ph of saturated solution mn(Oh )2 is 9.90 calculate oh ,mn, mnoh2 concentration and , ksp .compare calculated ksp value from the literature ksp value . are the two values close or different .
Answer #1

Answer:-

1. LiteratureKsp 1.9 X 10-13 and  Ksp = 2.5 x 10-13

both Ksp values are close to each other

2. Mn+1 = 3.97 x 10-5M

3. JOH-] = 7.94 x 10-5M

4. Mn(OH)2) = 3.97 10-M

Exlanation :-

The solubility equilibrium of Mn(OH)2 is

Mn(OH)2(6) = Mnt) + 20 Ha ...(1)

The Ksp of Mn(OH)2 is given as

Ksp = [Mn2+1[OH-] ....(2)

Let

the molar solubility of Mn(OH)2 in the solution be S (mol/L).

therefore,

Mn2+] = S(mol/L) and  [OH-] = 25(mol/L)

therefore, Ksp = S (25) ...(3)

We are given pH = 9.9

therefore,

pOH = 14 - pH = (14 - 9.9) = 4.1

Now,

p^{OH} = - log[OH^{-}]

ie. - log[OH^{-}] = 4.1

i.e log[OH^{-}] = - 4.1

ie. [OH^{-}] = antilog (- 4.1)

i.e JOH-] = 7.94 x 10-5M

ie. [OH^{-}] = 2S = 7.94 \times 10^{-5} M

from reaction (1) we have,

[Mn^{+}] = S = \frac{[OH^{-}]}{2} = \frac{7.94 \times 10^{-5} M}{2} = 3.97 \times 10^{-5} M

Therefore,

from eq(3) we have,

K_{sp} = 3.97 \times 10^{-5} \times (7.94 \times 10^{-5})^{2}

K_{sp} = 3.97 \times 10^{-5} \times 63.0436 \times 10^{-10}

K_{sp} = 2.50 \times 10^{-13}

LiteratureKsp 1.9 X 10-13

Now,

from, eq(3)

Ksp = S (25)  

Ksp = 4(S)

(S)^{3} = \frac{ K_{sp}}{4} = \frac{2.50 \times 10^{-13}}{4}

(S)^{3} = 6.25 \times 10^{-14}

taking cube root of both side we get,

S = 3.97 \times 10^{-5}

i.e [Mn(OH)_{2}] = 3.97 \times 10^{-5}